Find the work down in moving an object along a vector r = 3i + 2j - 5k,if the applied force is F = 2i-5 -k

5
4
9
3
1
Expert's answer
2015-02-27T10:18:06-0500
The work down in moving an object along a vector r, if the applied force F, is given by the dot product
A=Fr=(2i-5j-k)(3i+2j-5k)=2*3+(-5)*2+(-1)*(-5)=6-10+5=1.